# LLM-Craft: Robotic Crafting of Elasto-Plastic Objects with Large Language Models

### v2

@@ IV-B Single-Step Goals
Before we leverage the LLM for longer-horizon shape crafting goals, we first need to evaluate how well an LLM can out-of-the-box select the single action to achieve a single-step goal. While this is a relatively simple task, it proved somewhat difficult for the LLM due to how it is reasoning about the clay moulding task. We found that the LLM prescribes semantically meaningful descriptions for the clay states and goals in order to generate action sequences. This is a very useful attribute for the long horizon shape crafting task, but can interfere with performance for the single-step goals in which the goal shape isn’t anything in particular. For more detailed analysis of the LLM reasoning, please see the Supplemental Materials. We found the LLM was able to successfully reach the goal shape, but could sometimes get confused about rotation. The numerical results for the single-step experiments are shown in Table I, and the qualitative results are visualized in Figure 4.
Figure 5: Qualitative Shape Results. We compare the goal shapes to the performance of the LLM system without replanning, the full iterative proposed LLM-Craft pipeline, and a human demonstration baseline with the same discrete action representation.
@@ IV-C Long-Horizon Shape Goals
To further investigate the capabilities of LLMs for crafting, we evaluated the full system performance on a set of long-horizon letter shape goals consisting of the shapes C, I, L, T, X. We selected these letters because they are all able to be created with only deformation actions, whereas letters such as P are not because our current action formulation cannot create holes. To determine the effectiveness of the proposed iterative prompting scheme, we compared the performance with replanning enabled compared to executing the entire LLM-generated trajectory without replanning. Additionally, we compare the iterative LLM-Craft with an iterative variant in which the LLM also controls how much the end-effector is squeezed with options ’min’, ’medium’ and ’max’. Finally, to disambiguate the system errors from the limitations of the coarsely discretized action space, we conducted a set of human baseline experiments in which the human operator must select the two cells to squeeze together. The qualitative results are visualized in Figure 5, and the quantitative results are shown in Table II. Beyond presenting the CD/EMD similarity metrics with respect to the goal, we conduct human evaluation surveys and present the human classification accuracy for the shapes created by each experiment variant. For more details of how the human evaluation surveys were conducted, please see the appendix. We find that the LLM-Craft framework is able to successfully create the simple letter shaped sculptures at a comparable level to our human baseline with the discrete action space. Additionally, LLM-Craft can create smoother shapes when it also controls how much the gripper squeezes the clay. However, it is important to note that the performance of LLM-Craft has a high variance, as sometimes the system selects a bad action early into the trajectory and even with the iterative re-planning framework, the system is unable to recover.
COULD INCLUDE EXPERIMENTS FOR TEXT-ONLY GOALS!
TODO: get the squeeze variant human eval results
TODO: update the project website
Loop CD $\downarrow$ EMD $\downarrow$ Human Eval. $\uparrow$
[…]
Text Goal - - -
[…]
TABLE III: Quantitative Grid Ablation Results.
CD $\downarrow$ EMD $\downarrow$
[…]
Figure 6: Qualitative Grid Ablation Results. We compare the performance of LLM-Craft compared to a human baseline for variable grid sizes.
@@ IV-D Grid Ablation
To evaluate if the proposed framework can scale successful with different grid sizes, we deployed the LLM-Craft iterative pipeline on hardware for the ‘X’ letter task with grids of 6x6, 8x8, 10x10, 12x12 and 16x16. We conduct 5 experimental runs for each grid and present the visualization of the shape results in Figure 6 and the quantitative results in Table III. Additionally, we extended the human baseline in which a human selects actions with the same discrete action space to better visualize the shape capabilities of the constrained state/action space. For this ablation experiment, we kept the 3D printed grippers the same across experiments for the sake of time, meaning the varying grid size allowed for finer control of end-effector position, but the same quantity of clay was still being moved with each squeeze. Given the results of these experiments, we believe that the level of detail of shapes created could be more refined with the combination of these finer grids and thinner fingertips.
@@ IV-E Prompt Ablation
Figure 7: Prompt Ablation. The performance of the long-horizon, iterative system on the L shape task as we remove different components of the prompt. For full details of the prompt variations, please see the Supplemental Materials. For each prompt variation, we conducted 5 hardware experimental runs. We report the mean EMD and CD between the final shape and goal point clouds, with the black bar indicating the standard deviation across experiments.
To evaluate the impact of different components of the prompts on the system performance, we conducted an ablation study removing different parts. In particular, we chose to remove the following: (1) asking the LLM to predict the effect of the grasp on the clay after selecting an action, (2) asking the LLM to describe the similarities and differences between the state and goal before generating an action sequence, (3) removing the termination prompt altogether, (4) removing the goal text description, i.e. if the goal is an L to not mention L in the prompt explicitly, (5) removing the explanation of grasp types, specifically horizontal, vertical, or diagonal, (6) removing the instructions to choose action step-by-step, and (7) removing the explanations of the gridded state and action space. For the complete prompt and prompt variations for each ablation experiment, please see the Supplemental Materials. The breakdown of earth mover’s distance for the L shape across the ablations is shown in Figure 7. We chose to conduct the ablation experiments with the L shape goal because it is difficult to successfully create the bottom perpendicular line in the L given the discrete action space. Therefore, we believe this goal shape will help illuminate behavior differences given the changes to the prompt. Based on our ablation experiments, it is clear that removing the explanation of the grid from the prompt negatively impacts performance the most. This is an expected result, as the choice of state and action space in LLM-Craft is unique and requires a clear explanation, particularly of the labeling system of the grid in order to select the grid cells for actions. Removing the step-by-step instruction also significantly impacted performance, following the results from [kwon2023language]. We found that the grasp orientation description, goal text description and the termination prompt all had a mid-sized impact on the final performance of LLM-Craft. Finally, the requirements to reason about the state and goal as well as the predicted effect before selecting a grasp action had minimal impact on the average performance of LLM-Craft, but the variance was substantially reduced with their inclusion in the prompt. This makes sense, as requiring the LLM to reason about the differences between the current state and goal as well as describe physically how a grasp will change the state has the potential to reduce the frequency of physically unreasonable actions to be chosen. However, it is important to highlight that while all these components of the prompt reduce the unreasonable actions selected by the LLM, LLM-Craft does sometimes select actions and the explanation of the predicted effect is not fully physically accurate.
@@ IV-F Semantic Goals
Figure 8: Qualitative Semantic Goal Results. Given a semantic goal, i.e. how to qualitatively change the state without an explicit goal, LLM-Craft is able to successfully change the starting state to accordingly.
Up until this point, all of our experiments as well as past literature for elasto-plastic object robotic crafting has focused on explicit shape goals [bartsch2023sculptbot, bartsch2024sculptdiff, shi2022robocraft, bauer2024doughnet, shi2023robocook]. However, the core goal of creating a sculpture is to represent a shape or an idea, not not necessarily exactly replicate an existing sculpture or object. These more semantic goals are incredibly difficult to capture, quantify and evaluate, which is why to the best of our knowledge no prior work has explored this. However, given that we are using language in the LLM-Craft framework, we have a system that is more aligned with understanding and creating these semantic-based goals. For example, the termination prompt is evaluating more abstract shape similarity as opposed to low-level point similarity. We conducted a small set of experiments to evaluate if and how LLM-Craft handles more ambiguous semantic goals. We modified the action prompt to remove the inclusion of the goal image and description and instead included semantic commands such as ”make it thinner”, ”make it bumpy”, or ”make it shorter”. The qualitative results of these experiments are visualized in Figure 8. We found that the LLM-Craft framework was indeed able to modify the starting shape to better match these semantic goals. Even with the restricted action space, we are able to visually see the difference in how LLM-Craft chose to change the starting state to reach these semantically different goals.
@@ V Conclusion
In this work, we present LLM-Craft, to the best of our knowledge the first LLM-based robotic shape crafting system. Through real-world experiments, we demonstrate the powerful planning capabilities of LLMs out-of-the-box for both single-step and longer horizon shape goals. However, we find that while the semantic reasoning capabilities of LLMs are very powerful, LLM-based systems sometimes hallucinate bad actions. In the LLM-Craft system there is no real-world grounding, meaning sometimes the LLM selects actions and the explanation of the predicted effect of the action does not make any physical sense. However, we believe the semantic reasoning capabilities of LLMs could be very useful in future work combining the more classical and physics-based approaches with LLMs for the higher-level reasoning and semantic similarity analysis.

### v3

@@ IV-B Single-Step Goals
Before we leverage the LLM for longer-horizon shape crafting goals, we first need to evaluate how well an LLM can out-of-the-box select the single action to achieve a single-step goal. While this is a relatively simple task, it proved somewhat difficult for the LLM due to how it is reasoning about the clay molding task. We found that the LLM prescribes semantically meaningful descriptions for the clay states and goals in order to generate action sequences. This is a very useful attribute for the long horizon shape crafting task, but can interfere with performance for the single-step goals in which the goal shape isn’t anything in particular. We found the LLM was able to successfully reach the goal shape, but could sometimes get confused about rotation. The numerical results for the single-step experiments are shown in Table I, and the qualitative results are visualized in Figure 3. For these single-step experiments, we also explored an alternative state representation that is compatible with the grid-based action space. We implemented a baseline in which we represent the clay state as a binary array with a 1 if more than half of the cell contains clay and a 0 if less than half the cell contains clay. While both this and the gridded image representations provide sufficient information for the crafting task, we argue that the image-based representation allows for shape-level reasoning more easily. This is reflected in the experimental results in which the gridded image is visually able to more consistently match the goal shape. While the LLM can get confused about rotation with both representations, this happens less frequently with the gridded image. The image representation inherently has much more information than the simple binary array.
Figure 4: Qualitative Shape Results. We compare various LLM rollout strategies (no replanning, iterative, SC, ToT), the case of controlling the strength of the squeeze, semantic sculpting versus a ground truth goal image, and a human oracle.
Method Squeeze G. Img? CD [mm] $\downarrow$ EMD [mm] $\downarrow$ Curv. $\downarrow$ PAR $\downarrow$ H. Cls. $\uparrow$ H. Qual. $\uparrow$
Gemini 1.0 No Replan Fixed Yes 5.6 $\pm$ 1.9 8.2 $\pm$ 2.6 0.089 $\pm$ 0.012 0.047 $\pm$ 0.002 0.63 $\pm$ 0.17 4.4 $\pm$ 1.5
Iterative Fixed Yes 4.5 $\pm$ 1.5 6.7 $\pm$ 2.0 0.099 $\pm$ 0.021 0.048 $\pm$ 0.006 0.52 $\pm$ 0.34 5.9 $\pm$ 1.1
Iterative Varied Yes 5.5 $\pm$ 2.2 7.8 $\pm$ 2.5 0.088 $\pm$ 0.008 0.050 $\pm$ 0.006 0.68 $\pm$ 0.36 5.0 $\pm$ 1.5
Gemini 2.0 Iterative Varied Yes 3.8 $\pm$ 1.0 7.7 $\pm$ 2.4 0.102 $\pm$ 0.008 0.059 $\pm$ 0.010 0.81 $\pm$ 0.12 4.9 $\pm$ 1.5
Iterative Varied No - - 0.107 $\pm$ 0.011 0.055 $\pm$ 0.009 0.74 $\pm$ 0.17 4.3 $\pm$ 1.1
SC Varied Yes 3.7 $\pm$ 0.9 8.0 $\pm$ 1.9 0.095 $\pm$ 0.018 0.055 $\pm$ 0.007 0.83 $\pm$ 0.37 6.4 $\pm$ 1.3
SC Varied No - - 0.097 $\pm$ 0.005 0.059 $\pm$ 0.009 0.83 $\pm$ 0.37 5.3 $\pm$ 1.2
ToT Varied Yes 4.2 $\pm$ 1.1 8.5 $\pm$ 1.7 0.086 $\pm$ 0.014 0.059 $\pm$ 0.010 0.85 $\pm$ 0.20 5.4 $\pm$ 1.2
ToT Varied No - - 0.114 $\pm$ 0.017 0.063 $\pm$ 0.010 0.85 $\pm$ 0.34 6.3 $\pm$ 1.3
Human Robot Fixed Yes 5.1 $\pm$ 2.9 8.2 $\pm$ 3.7 0.090 $\pm$ 0.011 0.060 $\pm$ 0.010 0.80 $\pm$ 0.13 5.5 $\pm$ 1.3
Robot Varied Yes 2.6 $\pm$ 0.6 6.2 $\pm$ 1.2 0.110 $\pm$ 0.016 0.060 $\pm$ 0.010 1.00 $\pm$ 0.00 7.7 $\pm$ 1.0
TABLE III: API Cost Results. For each method with Gemini 2.0 flash across the shape goals of C, I, L, T, X.
Method G. Img? Cost [USD] $\downarrow$ # Actions
Iterative Yes 0.0037 $\pm$ 0.0019 5.16 $\pm$ 2.62
No 0.0039 $\pm$ 0.0013 5.80 $\pm$ 1.98
Self-Consistency Yes 0.0062 $\pm$ 0.0050 7.00 $\pm$ 4.18
No 0.0036 $\pm$ 0.0002 7.24 $\pm$ 1.99
Tree of Thought Yes 0.0113 $\pm$ 0.0007 7.60 $\pm$ 1.39
No 0.0097 $\pm$ 0.0014 6.48 $\pm$ 2.14
@@ IV-C Long-Horizon Shape Goals
To further investigate the capabilities of LLMs for crafting, we evaluated the full system performance on a set of long-horizon letter shape goals consisting of C, I, L, T, X. We selected these letters because they each require different skills and long-horizon strategies to achieve, and are all able to be created with only deformation actions, whereas letters such as P are not because our current action formulation cannot create holes. Through our experiments, we compare different rollout strategies (no replanning, iterative, self-consistency, and tree of thought), LLM and human action control with a fixed or varied gripper distance, and methods with and without a goal image provided. The fixed squeeze condition is one in which the robot always squeezes the end-effector to a fixed distance of 1.5cm between the fingertips. The varied squeeze condition allows the agent to select ’min’, ’medium’, or ’max’, corresponding to 2.5cm, 1.5cm, and 0.8cm respectively. The qualitative results are visualized in Figure 4, the quantitative results are shown in Table II, the LLM API costs and number of actions are shown in Table III, and a visualization of the human classification results across all methods are shown in Figure 5.
Through our experiments, we find that the LLM-Craft framework is able to successfully create visually identifiable shapes. In particular, we find that allowing the LLM to control both the gripper position and squeeze strength improves the quality of the final shapes created, particularly in terms of the human evaluations. Self-Consistency (SC) allowed for a selection of more consistent grasp strengths, which can be exemplified in the I shape results. Tree of thought (ToT) and SC rollout methods both outperform the iterative strategy in terms of shape smoothness, and human evaluation results. SC and ToT rollout strategies are able to more reliably generate nuanced sculpting trajectories, such as squeezing maximum strength on the first grasp and medium strength on the second perpendicular grasp to better define the final X shape without pushing the arms together. The improvement to shape quality by these sophisticated rollout methods can be best visualized for the X shape in Figure 4, and in improvements in human classification and quality ratings as compared to the iterative case. SC and ToT consistently are rated higher in terms of classification and quality (Welch’s t-test, $p_{SC}=0.06$, $t_{SC}=2.4$, $p_{ToT}=0.01$, $t_{ToT}=3.6$). Across the shapes, the human oracle with the ability to vary the squeeze of the gripper outperforms the best performing LLM-Craft variants of SC and ToT in quality rating by 1.3 and 2.3 respectfully (Welch’s t-test, $p_{SC}=0.006$, $t_{SC}=2.9$, $p_{ToT}=0.003$, $t_{ToT}=3.2$).
We find that CD/EMD are unreliable quantitative metrics for determining small differences in shape quality between our method variants. The CD/EMD metrics do not align well with the human evaluation results. This is because CD and EMD substantially penalize slight misalignment in terms of rotation, or shape structure with the specific goal point cloud, whereas the human raters prioritize shape recognizability as the goal letter and general visible smoothness. For the purpose of creating a semantically meaningful and recognizable shape, not all points on the surface are created equal. Certain deviations from the goal have little impact on human classification, whereas others have a significant impact. Some attributes of the shape are more important than others in terms of final classification and quality results. We hope to further explore and develop additional quantitative metrics that better align with human evaluations in future work.
The shape curvature and perimeter-to-area ratio are not quantitative metrics that alone can distinguish how well a created shape matches the goal. However, they do allow us to analyze nuanced differences between methods that perform similarly in terms of CD/EMD as well as human classification/quality ratings. In particular, while the SC method with and without a goal image have comparable human classification and quality rating results, the SC method with a goal image are ”smoother” in terms of lower scores in both the curvature and perimeter-to-area ratio across all shapes. Similarly, for ToT, the method with a goal image has lower curvature and perimeter-to-area ratio scores across all shapes. Both SC and ToT rollout methods consistently created letter shapes that were recognizable by human evaluators, with classification accuracy outperforming all other non-human methods (visualized in Figure 5). The primary shape these methods struggled to create, and human evaluators then struggled to classify was the C shape. This disparity between the C shape and the other letters is reflected in the larger standard deviations in the human evaluations. For a letter-by-letter breakdown of human evaluation results, please see the supplementary materials on the project website. Overall, the C shape was the most difficult shape for our system to create, largely due to the discrete action space. The C shape requires a smooth, tight curve which is very difficult to create with a discrete action grid. The challenge of this is reflected in the quality of C shapes the human oracle was able to create both with fixed/varied squeezes.
Figure 5: The confusion matrix of the human evaluations for all methods. A score of 1.0 indicates all human respondents classified the clay image as that letter.
TABLE IV: Quantitative Grid Ablation Results. We present the mean and standard deviation for the X shape with varying grids.
H. Qual. $\uparrow$ Curv. $\downarrow$ PAR $\downarrow$
6x6 Iterative 4.6 $\pm$ 2.6 0.059 $\pm$ 0.005 0.029 $\pm$ 0.003
Human 3.6 $\pm$ 2.5 0.074 $\pm$ 0.016 0.028 $\pm$ 0.001
8x8 Iterative 3.8 $\pm$ 3.1 0.064 $\pm$ 0.006 0.030 $\pm$ 0.001
Human 4.2 $\pm$ 2.7 0.070 $\pm$ 0.003 0.030 $\pm$ 0.003
10x10 Iterative 4.8 $\pm$ 2.4 0.068 $\pm$ 0.011 0.029 $\pm$ 0.003
Human 4.0 $\pm$ 2.9 0.073 $\pm$ 0.003 0.030 $\pm$ 0.003
12x12 Iterative 3.2 $\pm$ 3.3 0.077 $\pm$ 0.005 0.033 $\pm$ 0.003
Human 3.0 $\pm$ 2.3 0.074 $\pm$ 0.009 0.034 $\pm$ 0.002
16x16 Iterative 4.2 $\pm$ 2.9 0.085 $\pm$ 0.006 0.034 $\pm$ 0.001
Human 5.6 $\pm$ 2.4 0.081 $\pm$ 0.016 0.033 $\pm$ 0.004
@@ IV-D Grid Ablation
Figure 6: Qualitative Grid Ablation. We compare the performance of LLM-Craft compared to a human baseline for variable grid sizes.
To evaluate if the proposed framework can scale successfully with different grid sizes, we deployed the LLM-Craft iterative pipeline on hardware for the X letter task with grids of 6x6, 8x8, 10x10, 12x12 and 16x16. We conduct 5 experimental runs for each grid and present the visualization of the shape results in Figure 6 and the quantitative results in Table IV. Additionally, we extended the human baseline in which a human selects actions with the same discrete action space to better visualize the shape capabilities of the constrained state/action space. For this ablation experiment, we kept the 3D printed fingertips the same across experiments, meaning that varying the grid size allowed for finer control of the end-effector position, but the same quantity of clay was being moved with each squeeze. Through these experiments, we find that LLM-Craft can scale to grid sizes of 6x6, 8x8, and 10x10, with comparable quality, curvature and perimeter-to-area ratio mean values. However, when the grid gets particularly large (around 12x12 or 16x16), it becomes much more difficult to select quality actions, and LLM-Craft creates worse quality shapes.
@@ IV-E Prompt Ablation
To evaluate the impact of different components of the prompts on the performance, we conducted an ablation study removing different parts of the prompt. We conducted a second human evaluation survey for this ablation study in which we provided the evaluators with an image of the X shape created by each prompt variant. In this case, the evaluators knew that the shape was supposed to be an X, and were asked to rate the quality of the X shape on a scale of 1-10. In particular, we chose to remove the following (the order reflects the order shown in Figure 7): (1) goal image, (2) asking the LLM to describe the similarities and differences between the state and the goal, (3) asking the LLM to predict the effect of the grasp on the clay after selecting an action, (4) the explanation of grasp types, specifically horizontal, vertical, or diagonal, (5) the instructions to choose actions step-by-step, (6) the explanations of the gridded state and action space, (7) explaining the clay behavior when a cell is squeezed, (8) the goal text description (i.e. ’X’), and (9) the termination prompt. The breakdown of the human evaluations for the X shape across the ablations in shown in Figure 7.
Figure 7: Prompt Ablation. The performance of the long-horizon, iterative system on the X shape task as we remove different components of the prompt. For each prompt variation, we conducted 5 hardware experimental runs. We report the mean human quality evaluation, with the black bar indicating the standard deviation.
Based on our ablation experiments, the removal of the termination module negatively impacts performance the most. This is an expected result, because without the ability to recognize when the goal shape has been reached, the agent may continue to squeeze the clay and worsen the final shape created. The explanation of clay dynamics and the inclusion of the goal in text form have a large impact on the final shape quality. The explanation of clay dynamics is particularly important, as it explains how to conceptualize the changes in the cell-based state after a squeeze action. There was a medium impact on the final shape quality when the explanation of the grid and the step-by-step requirements were removed from the prompt. The reasoning about the predicted effect, grasp explanation, and the goal image had a small impact to the final shape quality. Finally, the requirement to reason about the state and goal had no effect on the final shape quality compared to the full prompt.
@@ IV-F Semantic Goals
In general, much of past literature for elaso-plastic object robotic crafting has focused on explicit shape goals [2, 4, 1, 5, 3]. However, the core goal of creating a sculpture is to represent a shape or an idea, not necessarily exactly replicate an existing sculpture or object. These more semantic goals are incredibly difficult to capture, quantify and evaluate. However, given that we are using language in the LLM-Craft framework, we have a system that is more aligned with understanding and creating these semantic-based goals. For example, the termination prompt is evaluating more abstract shape similarity as opposed to low-level point similarity. To evaluate how well our framework can handle these semantic goals, we conducted a variation of each rollout method in which we removed the goal image from the prompt. The results of these variants are shown in Figure 4 and Table II. We found that the LLM-Craft framework was indeed still able to create the simple letter shapes with a text-based goal alone. Without a goal image for guidance, the iterative framework would often stop earlier, whenever the termination agent would initially identify the letter in the clay. This resulted in an average of fewer actions, and final shapes that are recognized less often as compared to the iterative rollout with goal images (0.74 versus 0.81 mean classification accuracy, respectively). However, this discrepancy is resolved with the more sophisticated rollout strategies of SC (0.82 mean classification accuracy for both) and ToT (0.85 mean classification accuracy for both). Interestingly, the methods with text-based goals also have a wider variety of instances of the letters as compared to those with an explicit goal image. For example, the I shape created by Tree of Thought without a goal image includes the small horizontal top/bottom lines on the I shape that appear in some fonts. ToT without a goal image opts to squeeze the two most inward cells, leaving the small horizontal top/bottom lines on the I shape that appear in some fonts. While these are relatively simple semantically meaningful shapes, we hope to extend this exploration of semantic goals for 3D sculpture in future work.
@@ V Conclusion
In this work, we present LLM-Craft, to the best of our knowledge the first LLM-based robotic shape crafting system. Through real-world experiments, we demonstrate the powerful planning capabilities of LLMs out-of-the-box for both single-step and longer horizon shape goals. We find that LLMs contain relevant world knowledge for the clay crafting task without requiring any task-specific data, such as when creating an I one must squeeze along a single axis, versus when creating an X one must squeeze perpendicular to the previous grasp. For broader applications, our results demonstrate the usefulness of incorporating LLMs into robotic domains that typically require datasets that are task-specific, hardware-specific, and difficult to acquire. Additionally, we find that implementing more sophisticated rollout strategies such as Self-Consistency and Tree of Thought improves performance further and, in particular, reduces the frequency in which the system enters an irrecoverable clay state. We believe that the world knowledge and semantic reasoning capabilities of LLMs could be very useful in future deformable object manipulation work that combines the more classical and physics-based approaches with LLMs to reduce real-world dataset needs, and incorporate LLMs for higher-level reasoning and semantic similarity analysis.
Although our results demonstrate that LLM-Craft is effective in creating simple letter shapes, there is a clear limitation in shape complexity due to the grid-based action space. In particular, our framework struggles with the C shape because it is difficult to create a smooth, tight curve with a discrete, grid-based action space. We would expect similar limitations if LLM-Craft is extended to more complex shapes without developing a continuous action space. This framework could be extended to a continuous top-down action space by employing iterative visual prompting strategies similar to those in [33]. Furthermore, this framework could be extended to fully 3D shapes by incorporating multiple camera views, or depth images. We hope to explore these approaches in future work.